Considering the IQ results, the school administrator is also planning to offer additional math courses to the top 20%. What should be the IQ score of the student to be qualified to take the additional math course?


1
Expert's answer
2022-05-31T11:23:18-0400
"P(X>x)=1-P(Z\le\dfrac{x-108}{10})=0.2"

"P(Z\le\dfrac{x-108}{10})=0.8"

"\dfrac{x-108}{10}=0.8416"

"x=116.4"

The IQ score will be above 116.4